OpenStudy (ayonnaleflore):

Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden? 13 m wide 14 m wide 15 m wide 16 m wide

OpenStudy (error1603):

Welcome

OpenStudy (dinamix):

(x+10)(x+13)=208

OpenStudy (error1603):

(10+x)*(13+x) = 208 130+10x+13x+x^2 = 208 x^2+23x+130 = 208 x^2+23x+130-208 = 0 x^2+23x-78 = 0 (x-3)(x+26) = 0 x-3=0 or x+26 = 0 x = 3 or x = -26 Toss out the negative solution x = 3 is the only practical solution New Width: 10+x = 10+3 = 13 New Length: 13+x = 13+3 = 16

OpenStudy (error1603):

Hope it be helping,right?

OpenStudy (error1603):

Lol

OpenStudy (dinamix):

@Error1603 dont give her direct answer let think @ayonnaleflore

OpenStudy (ayonnaleflore):

yes it does ... thank you

OpenStudy (anonymous):

Sorry i couldn't help on this one i was helping someone else...

OpenStudy (error1603):

That's not a direct answer dude.